# Integrating GNN and Neural ODEs for Estimating Two-Body Interactions in Mixed-Species Collective Motion

### v1

@@ Abstract
Analyzing the motion of multiple biological agents, be it cells or individual animals, is pivotal for the understanding of complex collective behaviors. With the advent of advanced microscopy, detailed images of complex tissue formations involving multiple cell types have become more accessible in recent years. However, deciphering the underlying rules that govern cell movements is far from trivial. Here, we present a novel deep learning framework to estimate the underlying equations of motion from observed trajectories, a pivotal step in decoding such complex dynamics. Our framework integrates graph neural networks with neural differential equations, enabling effective prediction of two-body interactions based on the states of the interacting entities. We demonstrate the efficacy of our approach through two numerical experiments. First, we used a simulated data from a toy model to tune the hyperparameters. Based on the obtained hyperparameters, we then applied this approach to a more complex model that describes interacting cells of cellular slime molds. Our results show that the proposed method can accurately estimate the function of two-body interactions, thereby precisely replicating both individual and collective behaviors within these systems.
@@ 1 Introduction
Collective motion, a phenomenon observed in various biological systems, is characterized by the coordinated movement of multiple entities. This behavior is prevalent in a wide range of systems[26], from flocks of birds[5, 11, 7] and schools of fish[15] to cellular slime molds[12], microswimmers[10], and human crowds[17]. Understanding the underlying mechanisms of collective motion is crucial for elucidating the principles governing the dynamics of these systems. In particular, the interactions between individual entities play a pivotal role in shaping the collective behavior of the system.
Recent advances in imaging technologies have enabled detailed observations at the cellular level, providing insights into the dynamics of complex tissue formation involving multiple cell types. For example, cellular slime molds, a model organism for studying collective motion, exhibit intricate behaviors such as aggregation, migration, and differentiation[12]. These behaviors are driven by the interactions between different cell types, which are mediated by chemical signals and physical forces. Decoding the underlying equations of motion governing these interactions is essential for understanding the emergent properties of these systems.
In this work, we present a novel deep learning framework for estimating two-body interactions in mixed-species collective motion. Our framework integrates graph neural networks (GNNs) with neural differential equations (neural DEs) to predict the interactions between pairs of entities based on their states. GNNs are well-suited for modeling complex interactions in graph-structured data, while neural DEs provide a flexible framework for learning the dynamics of the system. By combining these two approaches, we can effectively capture the interactions between individual entities and predict their collective behavior. We demonstrate the efficacy of our framework through two numerical experiments. The first experiment uses a toy model designed to generate data for refining the hyperparameters of our framework. The second experiment explores a complex scenario mimicking the collective motion of cellular slime molds, where two different cell types interact with each other. Our results show that our method can accurately estimate the two-body interactions, thereby replicating both individual and collective behaviors within these systems.
The rest of this paper is organized as follows. In Section 2, we introduce the study of collective motion. In section 3, we provide an overview of related work on collective motion and deep learning for dynamical systems. In Section 4, we describe our deep learning framework for estimating two-body interactions in mixed-species collective motion. In Section 5, we present the results of our numerical experiments. Finally, in Section 6, we discuss the implications of our work and outline future research directions.
@@ 2 Background
Let us first describe the formulation of collective motion in active matter. Starting with the Vicsek model [25], the collective behavior in active matter is described based on the centroids, velocities, and orientations of each individual. A unique aspect of active matter is that it allows for the spontaneous generation of forces and torques, which is justified by the ability to extract and utilize energy from the external environment[21]. The Vicsek model itself is a multi-particle model in discrete time, where each individual possesses a velocity along its orientation and adjusts its direction based on interactions with nearby individuals. This model assumes that interactions among individuals are local and can be represented using a graph structure[1]. Stemming from this model, numerous other models have been proposed, differing in the nature of interactions between individuals and the forms of their equations of motion. Particularly models with continuous-time motion equations often use Langevin equations where the motion of individuals is described by the summation of pairwise interactions[8]. Furthermore, since active matter can utilize energy from the external environment, focusing solely on the moving agents categorizes it as an open system. This implies that the conservation of total system energy is not required, making these systems inherently unsuitable for Hamiltonian descriptions. Given this background, the formulation of collective motion in active matter typically involves direct descriptions of forces rather than using free energy.
@@ 3 Related Work
This section discusses methods for estimating governing laws from data, commonly referred to as system identification. System identification aims to estimate the equations of motion of a system from data. One well-known method for estimating differential equations is Sparse Identification of Nonlinear Dynamics (SINDy) [6]. SINDy estimates nonlinear differential equations from data using LASSO regression to sparsely estimate the terms of the equations. However, when applying this method to many-body systems with pairwise interactions, the number of parameters to estimate increases exponentially with the number of individuals, which raises computational costs and leads to potential instabilities. Other system identification methods use Bayesian optimization [3, 4, 2], e.g. representing the functions in the motion equations through basis function expansion, and estimating their coefficients via Bayesian optimization. This method has been successfully used in Vicsek model, reproducing its formation of orientational order. However, the effectiveness of this approach for more complex systems, such as mixed-species systems, remains unclear. Deep learning approaches to analyzing collective motion have also been proposed [13, 24, 18]. These incorporate Attention mechanisms to analyze behaviors limited to turning right or left [13], systems that limit the scope to Vicsek-type models to estimate orientational order parameters [24], or to estimate two-body interaction as a function of and those using Hamiltonians to estimate repulsive forces [18]. Nevertheless, these methods have limitations regarding the types of systems they can be applied to or the specific behaviors they can estimate. Particularly, no methods have been proposed so far that can be applied to systems where multiple species interact in unknown ways.
@@ 4 Method
The goal of this study is to estimate the rules of motion for individual entities within collective motion data. We represent the state of each entity $i$ at time $t$ as $\mathcal{Z}\ni z^{i}(t)=(x^{i}(t),y^{i})\in\mathcal{X}\times\mathcal{Y}$, where $x^{i}\in\mathcal{X}$ denotes the variables described by the motion equations in the state space $\mathcal{X}\subset\mathbb{R}^{n}$, and $y^{i}$, a non-temporal variable, represents auxiliary attributes such as the type of each entity within the feature space $\mathcal{Y}$. We define a distance function $d:\mathcal{Z}\times\mathcal{Z}\to\mathbb{R}$, and assume that entities $i$ and $j$ interact if $d(z^{i}(t),z^{j}(t))<d_{0}$, where $d_{0}\in\mathbb{R}$ is a predefined threshold. Consequently, the system at any time $t$ can be represented as an undirected graph $G(t)=(V,E(t),Z(t))$, where $V=\{1,2,\ldots,N\}$ is the set of entities, $E(t)\subset V\times V$ denotes the pairs of interacting entities, and $Z(t)=\{z^{i}(t)|i\in V\}$ is the set of states of all entities. The motion of each entity $i\in V$ is governed by the following Langevin equation:
$dx^{i}=\left(F^{(1)}(z^{i}(t))+\sum_{j\,\text{s.t.}\,(i,j)\in E(t)}F^{(2)}(z^{i}(t),z^{j}(t))\right)dt+\sigma dW^{i}(t),$ (1)
where $F^{(1)}:\mathcal{Z}\to\mathbb{R}^{n}$ represents self-driven forces, $F^{(2)}:\mathcal{Z}\times\mathcal{Z}\to\mathbb{R}^{n}$ encapsulates the forces due to interactions between pairs, $\sigma\in\mathbb{R}$ is the intensity of noise, and $W^{i}(t)\in\mathbb{R}^{n}$ is a Wiener process.
Given the initial state $Z(t_{0})\in\mathcal{Z}$ at time $t_{0}\in\mathbb{R}$, the state $Z(t_{0}+\Delta t)\in\mathcal{Z}$ after a time interval $\Delta t\in\mathbb{R}$ can be determined by solving the motion equation using the operator $S(F^{(1)},F^{(2)},\sigma,d,d_{0}):(Z(t_{0}),\Delta t)\mapsto Z(t_{0}+\Delta t)$. Here, $z^{i}(t_{0}+\Delta t)=(x^{i}(t_{0}+\Delta t),y^{i})$ is defined by the following integral:
@@
which is computed numerically. For $\sigma=0$, we use the neural ordinary differential equation (neural ODE) method[9, 23], and for $\sigma\neq 0$ the neural stochastic differential equation (neural SDE) method[19, 16]. The summation over $E(t)$ in the equations is efficiently calculated using a graph neural network (GNN) based on the message passing algorithm. To facilitate these computations within the neural DE frameworks, we developed a wrapper using a GNN framework, deep graph library[27]. This wrapper computes the set $E(t)$ using the provided $z(t)$, $d$, and $d_{0}$, and then calculates and returns the summation of forces $F^{(1)}$ and $F^{(2)}$. In particular, this constitutes a concrete implementation of $S(F^{(1)},F^{(2)},\sigma,d,d_{0})$.
4.1 Simulation with a Given Model
Initially, we used the developed wrapper to conduct simulations with a predefined model. Specific forms for the interaction and distance functions $F^{(1)}_{\text{sim}}$, $F^{(2)}_{\text{sim}}$, and $d_{\text{sim}}$ were provided, and appropriate settings for the noise magnitude $\sigma_{\text{sim}}$ and threshold $d_{0,\text{sim}}$ were configured to construct $S_{\text{sim}}=S(F^{(1)}_{\text{sim}},F^{(2)}_{\text{sim}},\sigma_{\text{sim}},d_{\text{sim}},d_{0,\text{sim}})$. For each $i\in V$, initial states $x^{i}(0)\in\mathcal{X}_{0}$ were assigned uniformly at random, and appropriate values for $y^{i}\in\mathcal{Y}$ were also specified. Using $S_{\text{sim}}$, the motion equations were solved. The simulation was repeated $M$ times, and the $m$-th result was denoted as $Z^{m}_{\text{sim}}:t\mapsto S_{\text{sim}}(Z^{m}_{\text{sim}}(0),t)$ with the $m$-th initial condition $Z^{m}_{\text{sim}}(0)$. The combined data set $Z_{\text{sim}}=\{Z^{m}_{\text{sim}}|m=1,\ldots,M\}$ was then used to apply the subsequent learning algorithm, aiming to estimate $F^{(1)}_{\text{sim}}$ and $F^{(2)}_{\text{sim}}$.
@@ 4.1 Simulation with a Given Model
4.2 Learning Algorithm
Subsequently, we developed a learning algorithm to estimate $F^{(1)}_{\text{sim}}$ and $F^{(2)}_{\text{sim}}$ using the results from the simulations conducted. The self-propulsion and interaction functions, $F^{(1)}_{\text{NN}}$ and $F^{(2)}_{\text{NN}}$, were modeled using neural networks. Unless otherwise specified, these functions consisted of a three-layer fully connected network followed by a scaling layer. Each fully connected layer consisted of 128 nodes, with the Exponential Linear Unit (ELU) serving as the activation function. The scaling layer involved a scalar $A$ and a vector $B$ as parameters, and it transformed an input vector $\alpha\in\mathbb{R}^{n}$ to $e^{A}\alpha+B$. This configuration allowed the output of the fully connected network to be scaled appropriately, reflecting physical scales.
The parameters of the fully connected networks and the scaling layer, collectively denoted as $\theta$, were optimized to approximate $F^{(1)}_{\text{NN}}(\theta)$ and $F^{(2)}_{\text{NN}}(\theta)$ to $F^{(1)}_{\text{sim}}$ and $F^{(2)}_{\text{sim}}$, respectively. To evaluate the deviation of $F^{(1)}_{\text{NN}}(\theta)$ and $F^{(2)}_{\text{NN}}(\theta)$, we first solved the motion equations using these functions. Specifically, we constructed $S_{\text{NN}}(\theta)=S(F^{(1)}_{\text{NN}}(\theta),F^{(2)}_{\text{NN}}(\theta),0,d_{\text{sim}},d_{0,\text{NN}})$ and used it to perform simulations. Here, $d_{0,\text{NN}}$ was a suitably set threshold, and for simplicity, $\sigma_{\text{NN}}=0$. In the simulation, the initial state was set as $Z^{m}_{\text{sim}}(t_{0})$, and we used this to compute the state $\tau$ time units later as $Z^{m}_{\text{NN}}(\theta;t_{0}+\tau)=S_{\text{NN}}(\theta)(Z^{m}_{\text{sim}}(t_{0}),\tau)$.
A loss function $\mathcal{L}:\mathcal{Z}\times\mathcal{Z}\to\mathbb{R}$ was used to evaluate the discrepancy between the solution by $\theta$ and the simulation data through $\mathcal{L}^{m}(\theta;t_{0})=\mathcal{L}(Z^{m}_{\text{sim}}(t_{0}+\tau),Z^{m}_{\text{NN}}(\theta;t_{0}+\tau))$. To optimize $\theta$, the average $\mathcal{L}(\theta)$ of this loss function over various $m$ and $t_{0}$ was minimized:

### v2

# Integrating GNN and Neural ODEs for Estimating Non-Reciprocal Two-Body Interactions in Mixed-Species Collective Motion

@@ Abstract
Analyzing the motion of multiple biological agents, be it cells or individual animals, is pivotal for the understanding of complex collective behaviors. With the advent of advanced microscopy, detailed images of complex tissue formations involving multiple cell types have become more accessible in recent years. However, deciphering the underlying rules that govern cell movements is far from trivial. Here, we present a novel deep learning framework for estimating the underlying equations of motion from observed trajectories, a pivotal step in decoding such complex dynamics. Our framework integrates graph neural networks with neural differential equations, enabling effective prediction of two-body interactions based on the states of the interacting entities. We demonstrate the efficacy of our approach through two numerical experiments. First, we used simulated data from a toy model to tune the hyperparameters. Based on the obtained hyperparameters, we then applied this approach to a more complex model with non-reciprocal forces that mimic the collective dynamics of the cells of slime molds. Our results show that the proposed method can accurately estimate the functional forms of two-body interactions – even when they are nonreciprocal – thereby precisely replicating both individual and collective behaviors within these systems.
@@ 1 Introduction
Collective motion, a phenomenon observed in various biological systems, is characterized by the coordinated movement of multiple entities. This behavior is prevalent in a wide range of self-propelled systems, collectively referred to as active matter [34], from flocks of birds [5, 13, 7] and schools of fish [19] to cellular slime molds [14], microswimmers [12], swarming leukocytes [32], and human crowds [21]. Understanding the underlying mechanisms of collective motion is crucial for elucidating the principles governing the dynamics of these systems. In particular, the interactions between individual entities play a pivotal role in shaping the collective behavior of the system.
Recent advances in imaging technologies have enabled detailed observations at the cellular level, providing insights into the dynamics of complex tissue formation involving multiple cell types. For example, cellular slime molds, a model organism for studying collective motion, exhibit intricate behaviors such as aggregation, migration, and differentiation [14]. These behaviors are driven by the interactions between different cell types, which are mediated by chemical signals and physical forces. Decoding the underlying equations of motion that govern these interactions is essential for understanding the emergent properties of these systems.
In this work, we present a novel deep learning framework for estimating two-body interactions in a mixed-species collective motion. Our framework integrates graph neural networks (GNNs) with neural differential equations (neural DEs) to predict interactions between pairs of entities based on their states. GNNs are well-suited for modeling complex interactions in graph-structured data, while neural DEs provide a flexible framework for learning the dynamics of the system. By combining these two approaches, we can effectively capture the interactions between individual entities and predict their collective behavior. We demonstrate the efficacy of our framework through two numerical experiments. The first experiment uses a toy model designed to generate data for refining the hyperparameters of our framework. The second experiment explores a complex scenario that partially mimics the collective motion of cellular slime molds, where two different cell types interact with each other. Our results show that our method can accurately estimate the two-body interactions, thereby replicating both individual and collective behaviors within these systems.
The rest of this paper is organized as follows. In section 2, we introduce the study of collective motion. In section 3, we provide an overview of related work on collective motion and deep learning for dynamical systems. In section 4, we describe our deep learning framework for estimating two-body interactions in mixed-species collective motion. In section 5, we successfully apply this framework to two numerical experiments. Finally, in section 6, we discuss the implications of our work and outline potential future research directions.
@@ 2 Background
Let us first describe the formulation of collective motion in active matter. Starting with the Vicsek model [33], the collective behavior in active matter is described based on the centroids, velocities, and orientations of each individual. A unique aspect of active matter is that it allows for spontaneous generation of forces and torques, which is justified by the ability to extract and utilize energy from the external environment [26]. The Vicsek model itself is a multi-particle model in discrete time, where each individual possesses a velocity along its orientation and adjusts its direction based on interactions with nearby individuals. This model assumes that interactions among individuals are local and can be represented using a dynamical graph structure [1]. Stemming from this model, numerous other models have been proposed, differing in the nature of interactions between individuals and the forms of their equations of motion. Particularly models with continuous-time motion equations often use Langevin equations where the motion of individuals is described by the summation of pairwise interactions [8]. Furthermore, since active matter can utilize energy from the external environment, focusing solely on moving agents categorizes it as an open system. This implies that the conservation of the total system energy is not guaranteed, making these systems inherently unsuitable for Hamiltonian descriptions. Given this background, the formulation of collective motion in active matter typically involves direct descriptions of forces rather than using free energy.
@@ 3 Related Work
System Identification methods:
This section discusses methods for estimating governing laws from data, commonly referred to as system identification. System identification aims to estimate the equations of motion of a system from data. One well-known method for estimating differential equations is Sparse Identification of Nonlinear Dynamics (SINDy) [6]. SINDy estimates nonlinear differential equations from data using LASSO regression to sparsely estimate the terms of the equations. However, when applying this method to many-body systems with pairwise interactions, the number of parameters to estimate increases exponentially with the number of individuals, which raises computational costs and leads to potential instabilities. Other system identification methods use Bayesian optimization [3, 4, 2], e.g. representing the functions in the motion equations through basis function expansion, and estimating their coefficients via Bayesian optimization. This method has been successfully used in Vicsek model, reproducing its formation of orientational order. However, the effectiveness of this approach for more complex systems, such as mixed-species systems, remains unclear. Deep learning approaches to analyzing collective motion have also been proposed [17, 31, 22]. These incorporate Attention mechanisms to analyze behaviors limited to turning right or left [17], systems that limit the scope to Vicsek-type models to estimate orientational order parameters [31], or to estimate two-body interactions by expressing forces as neural networks [22]. Another possibility is to express Hamiltonians [15] or Lagrangians [11] as neural networks, in order to preserve known symmetries of the studied system. Some proposed including knowledge or assumptions about the physical context into the loss function [25]. So far, none of these methods have been applied to systems in which multiple species interact with each other in unknown ways.
Integration of graph neural network with neural differential equations:
In this paper, we integrate GNN with neural ordinary differential equations (neural ODEs) to estimate interactions through evaluating equations of motion with dynamic edge structure. Most of the previous studies on integration of the two methods solve and train ODEs on a graph with fixed edges and consider changes in the edge weight [36, 29, 9, 16, 38]. This approach requires a fully or almost fully connected graph for collective motion where the adjacency changes drastically over time, and causes memory inefficiency. On the other hand, some introduced a given time series of edge structure [28], making the system difficult to be extrapolated. Thus we resolved these issues by setting a rule to define edges at each time point instead of explicitly giving graph structure time series.
@@ 4 Method
We aim to estimate the rules of motion for individual entities within collective motion data. This section presents the general framework that is used to generate the training data from a physical model for two numerical experiments, as well as the learning algorithm and neural network architecture.
4.1 Physical model
We represent the state of each entity $i$ at time $t$ as $\mathcal{Z}\ni z^{i}(t)=(x^{i}(t),y^{i})\in\mathcal{X}\times\mathcal{Y}$, where $x^{i}\in\mathcal{X}$ denotes the variables described by the motion equations in the state space $\mathcal{X}\subset\mathbb{R}^{n}$, and $y^{i}$, a non-temporal variable, represents auxiliary attributes such as the type of each entity within the feature space $\mathcal{Y}$. We define a distance function $d:\mathcal{Z}\times\mathcal{Z}\to\mathbb{R}$, and assume that entities $i$ and $j$ interact at time $t$ if $d_{ij}(t):=d(z^{i}(t),z^{j}(t))<d_{0}$, where $d_{0}\in\mathbb{R}$ is a predefined threshold. The motion of each entity $i$ is governed by the following Langevin equation:
$dx^{i}=\left(F^{(1)}(z^{i}(t))+\sum_{j\,\text{s.t.}\,d_{ij}(t)<d_{0}\wedge j\neq i}F^{(2)}(z^{i}(t),z^{j}(t))\right)dt+\sigma dW^{i}(t),$ (1)
with given self-driven forces $F^{(1)}:\mathcal{Z}\to\mathbb{R}^{n}$, forces due to interactions between pairs $F^{(2)}:\mathcal{Z}\times\mathcal{Z}\to\mathbb{R}^{n}$, intensity of noise $\sigma\in\mathbb{R}$, and a Wiener process $W^{i}(t)\in\mathbb{R}^{n}$. Given the initial state $Z(t_{0})\in\mathcal{Z}$ at time $t_{0}\in\mathbb{R}$, the state $Z(t_{0}+\Delta t)\in\mathcal{Z}$ after a time interval $\Delta t\in\mathbb{R}$ can be determined by solving the motion equation using the operator $S(F^{(1)},F^{(2)},\sigma,d,d_{0}):(Z(t_{0}),\Delta t)\mapsto Z(t_{0}+\Delta t)$. Here, $z^{i}(t_{0}+\Delta t)=(x^{i}(t_{0}+\Delta t),y^{i})$ is defined by the following integral:
@@ 4.1 Physical model
which is computed numerically (section 4.3). For $\sigma=0$, we evaluated eq.2 with the neural ODE method [10, 30], and for $\sigma\neq 0$ the neural stochastic differential equation (neural SDE) method [23, 20].
4.2 Interpretation of the system as a graph
The system at any time $t$ can be represented as a directed graph $G(t)=(V,E(t),Z(t))$, where the set of entities represent the vertices $V=\{1,2,\ldots,N\}$, pairs of interacting entities constitute edges $E(t)=\{(i,j)\in V\times V|d_{ij}(t)<d_{0}\wedge i\neq j\}$, and $Z(t)=\{z^{i}(t)|i\in V\}$ is the set of states of all entities. Importantly, since the entities change states over time, the graph in general evolves over time. The forces $F^{(1)}$ and $F^{(2)}$ are evaluated on each vertex and each edge, respectively.
Figure 1: Schematic figure of the integration of GNN and neuralODE. For details, see section 4.3.
4.3 Solving the equations of motion of the graph through neural DEs
To obtain a time series of collective entities (eq. 2), we have to evaluate the equation of motion (eq. 1). We achieved this with a custom-made wrapper that connects neural ODE (SDE) to GNN (Figure 1). The wrapper is equipped with the distance function $d$ and the threshold $d_{0}$ to define the edges $E(t)$ at each time $t$, the functions to calculate self-propulsion $F^{(1)}$ and interaction $F^{(2)}$, and the noise intensity $\sigma$. The computation $S(F^{(1)},F^{(2)},\sigma,d,d_{0})$ is then run through the following steps:
  1. 0.
    Create a graph object only with nodes $V$ and set the static states $\{y^{i}\}$ of the entities. Store this graph in the wrapper object. Pass the initial state $\{x^{i}(t_{0})\}$ into the neural ODE (SDE) object (Figure 1 blue).
  2. 1.
    The neural ODE (SDE) object passes the current state $\{x^{i}(t)\}$ to the forward function of the wrapper object.
  3. 2.
    The forward function places the state values $\{x^{i}(t)\}$ at the nodes of the graph. Subsequently, the graph is passed to GNN (Figure 1 red).
  4. 3.
    GNN first defines and updates the edges $E(t)$ using the distance function $d$ and the threshold $d_{0}$ stored in the wrapper. Next, it calculates the interactions $F^{(2)}$ on the edges and adds them to the self-propulsion $F^{(1)}$ on the nodes to return the forces to the ODE (SDE) object. If SDE, GNN also returns the noise intensity $\sigma$ in the wrapper.
  5. 4.
    The neural ODE (SDE) object updates the state $\{x^{i}(t+dt)\}$.
  6. 5.
    Repeat steps 1-4 until the end of the time series.
When inferring the forces $F^{(1)},F^{(2)}$ , we model them using neural networks (section 4.5).
4.4 Generation of training data
To construct $S_{\text{sim}}=S(F^{(1)}_{\text{sim}},F^{(2)}_{\text{sim}},\sigma_{\text{sim}},d_{\text{sim}},d_{0,\text{sim}})$, distance functions $d_{\text{sim}}$ were provided, suitable choices for the noise magnitude $\sigma_{\text{sim}}$, threshold $d_{0,\text{sim}}$ and forms for the forces $F^{(1)}_{\text{sim}}$, $F^{(2)}_{\text{sim}}$ were specified (see section 5). Thus defined $S_{\text{sim}}$ was used to conduct simulations according to the motion equations, from initial states assigned uniformly at random $x^{i}(0)\in\mathcal{X}_{0}$ , with appropriate values for $y^{i}\in\mathcal{Y}$. The simulation was repeated $M$ times, and the $m$-th result is denoted as $Z^{m}_{\text{sim}}:t\mapsto S_{\text{sim}}(Z^{m}_{\text{sim}}(0),t)$ with the $m$-th initial condition $Z^{m}_{\text{sim}}(0)$. The combined data set $Z_{\text{sim}}=\{Z^{m}_{\text{sim}}|m=1,\ldots,M\}$ was then used to apply the subsequent learning algorithm, aiming to estimate $F^{(1)}_{\text{sim}}$ and $F^{(2)}_{\text{sim}}$.
@@ 4.4 Generation of training data
4.5 Learning algorithm and neural network architecture for the forces
Subsequently, we developed a learning algorithm to estimate $F^{(1)}_{\text{sim}}$ and $F^{(2)}_{\text{sim}}$ using the results from the simulations conducted.
The self-propulsion and interaction functions, $F^{(1)}_{\text{NN}}$ and $F^{(2)}_{\text{NN}}$, were modeled using neural networks. Unless otherwise specified, these functions consisted of a three-layer fully connected network followed by a scaling layer. Each fully connected layer consisted of 128 nodes, with the Exponential Linear Unit (ELU) serving as the activation function. The scaling layer involved a scalar $A\in\mathbb{R}$ and a vector $B\in\mathbb{R}^{n}$ as learnable parameters, and it transformed an input vector $\alpha\in\mathbb{R}^{n}$ to $e^{A}\alpha+B$. This configuration allowed the output of the fully connected network to be scaled appropriately, reflecting physical scales.
The parameters of fully connected networks and the scaling layer, collectively denoted $\theta$, were optimized to approximate $F^{(1)}_{\text{NN}}(\theta)$ and $F^{(2)}_{\text{NN}}(\theta)$ to $F^{(1)}_{\text{sim}}$ and $F^{(2)}_{\text{sim}}$, respectively. To evaluate the deviation of $F^{(1)}_{\text{NN}}(\theta)$ and $F^{(2)}_{\text{NN}}(\theta)$, we first solved the motion equations using these functions. In practice, we constructed a wrapper object $S_{\text{NN}}(\theta)=S(F^{(1)}_{\text{NN}}(\theta),F^{(2)}_{\text{NN}}(\theta),0,d_{\text{sim}},d_{0,\text{NN}})$ and used it to perform simulations. Here, $d_{0,\text{NN}}$ was a suitably set threshold, and for simplicity, $\sigma_{\text{NN}}=0$. In the simulation, the initial state was set as a snapshot in training data $Z^{m}_{\text{sim}}(t_{0})$, and we used this to compute the state $\tau$ time units later as $Z^{m}_{\text{NN}}(\theta;t_{0}+\tau)=S_{\text{NN}}(\theta)(Z^{m}_{\text{sim}}(t_{0}),\tau)$, which would be compared with a truth $Z^{m}_{\text{sim}}(t_{0}+\tau)$.
For each numerical experiment, a loss function $\mathcal{L}:\mathcal{Z}\times\mathcal{Z}\to\mathbb{R}$ was used to evaluate the discrepancy between the solution by $\theta$ and the simulation data through $\mathcal{L}^{m}(\theta;t_{0})=\mathcal{L}(Z^{m}_{\text{sim}}(t_{0}+\tau),Z^{m}_{\text{NN}}(\theta;t_{0}+\tau))$. To optimize $\theta$, the average $\mathcal{L}(\theta)$ of this loss function over various $m$ and $t_{0}$ was minimized: